2) For the function y = x^2 - 4x - 5, perform the following tasks:
a) Put the function in the form y = a(x - h)2 + k.
Answer:
Show work in this space
b)What is the line of symmetry?
Answer:
c) Graph the function using the equation in part a. Explain why it is not necessary to plot points to graph when using y = a (x – h)2 + k.
Show graph here.
Explanation of graphing.
d) In your own words, describe how this graph compares to the graph of y = x^2?
Answer:

Answer by stanbon(48554)

a) y=(x^2-4x )-5
Complete the square:
y=(x^2-4x+4)-5-4
y=(x-2)^2-9
b)The line of symmetry is x=2
c) Draw the line of symmetry.
When x=2, y=-9; plot the point (2,-9)
That is the vertex of the parabola.
The y=intercept is (0,-5); plot it.
Draw the parabola which opens up.
d) The (x-2) moves y=x^2 two units to the right.
The "-9" moves the parabola nine units down.
Stan H.